OpenStudy (finn):

Solve the system of linear equations by substitution WILL GIVE MEDAL AND FAN! y = 5x + 6 can anyone provide an example of how to do it with the same set-up as my question above? something like.... y = 3x + 4 or something?

hartnn (hartnn):

i don't see a system of equations, i only see 1 equation y=5x+6 and yes.

hartnn (hartnn):

by substitution means you need to put y=.... from one equation into other equation.

hartnn (hartnn):

example, y= ax+b, y=cx+d then ax+b = cx+d and you can find 'x' from here.

hartnn (hartnn):

yes,

hartnn (hartnn):

5x + 6 = -2x - 8 add 2x on both sides, subtract 6 from both sides... in short, try to find 'x'

hartnn (hartnn):

where did 'x' go ? add "2x" on both sides.

hartnn (hartnn):

 yes, now simplify that.